The potential energy function associated with a force acting on a system is v - 3xy - 8x. What is the force at point (x,y)?
Answer #1

Sol nor potential Energy function U= 300g - 8u fora, F=-vu + dz 1-800 dn ay :-( (21x54-8) î+ 3x? to) F = (8-21 wy) î - 3x? 71